# Evidence of 1:1 slope between rocky Super-Earths and their host stars

### v1

@@ 1 Introduction
Starting with planetesimals, that are presumed to have homogenous composition (similar to the refractory ratios of the host star Larimer1967; Grossman1972) under efficient radial mixing in the protoplanetary disk, the planet grows through a series of processes that may or may not alter their primordial composition (Asphaug2006; Carter2015; Scora2020; Clement2021). In the Solar System, there is a moderate variety in major element composition as seen by the iron enrichment of Mercury and stony-iron meteorites as well as the iron-depletion of the Moon and aubrite meteorites, with values in between (Keil1989; Morgan1980; Lodders2003; Wieczorek2006; Margot2018). Proposed mechanisms that result in major chemical sorting so far include those that are based on differing Fe, Si and Mg properties early in the evolution, such as, fractionation driven by the differences in the condensation temperatures of Fe vs Mg-Si (Lewis1972; Grossman1972; Aguichine2020); photophoresis, which causes Fe vs Mg-Si dust grains to drift differently due to differences in conductivity and density when heated by stellar radiation (Krauss2005; Wurm2013; Cuello2016); increased nucleation of Fe vs Mg-Si pebbles due to differences in grain size and surface tension that affect growth under the streaming instability (Johansen2022); or those that are based on losing mantle material once the planet has differentiated into a silicate mantle over an iron core, such as giant impact collisions (Benz1988; Benz2007; Asphaug2014; Clement2021; Scora2024; Ferich2025) and for small bodies (i.e. Mercury-sized), photoevaporation of the mantle (Ito2021). Thus, comparing the relative abundances of planetary bodies to that of their star, offers the possibility to understand which mechanisms form planets.
@@
Chemical comparisons had been limited only to the Solar System or a few planets (Santos2015) until recently, when enough data of rocky exoplanets with good precision had grown to several tens of planets. Plotnykov2020 compared a population of allegedly rocky planets to that of the population of planet-hosting stars and found that planets exhibit a greater diversity in core-mass fraction (CMF), iron to magnesium (Fe/Mg) and iron to silicate (Fe/Si) ratios relative to the stellar population. Subsequent studies made one-to-one comparisons in an effort to uncover any possible correlations (Adibekyan2021; Schulze2021; Wang2022; Liu2023; Unterborn2023; Brinkman2024; Adibekyan2024). Notably, Adibekyan2021 suggested that super-Mercuries form a distinct population around stars that are enriched in core to mantle material (Fe/Mg+Si) and proposed a strong linear relationship with a slope steeper than a 1:1 line between iron-to-silicate mass fraction of planets and their host stars (slope of ). This result has been revised in Brinkman2024 and Adibekyan2024 works such that a super-Mercury subpopulation is no longer present, but the steep compositional slope remains.
@@
However, most of these different studies have a shortcoming that stems from considering separately the elemental abundances of stars versus their stellar mass and radius. Given that the latter affect inferences on the planetary mass-radius (M-R) and consequently the planetary interior composition, ensuring consistency between stellar composition, mass and radii is expected to update the planetary characteristics. To address this issue and ensure consistency, two groups (Brinkman2024; paperI) have reanalyzed the stellar data and updated planetary M-R accordingly. In our companion paper, hereafter paperI (paperI), we explain in detail the procedure to obtain a consistent stellar data set and perform simple planet-stars comparisons.
@@
In this paper, we delve deeper into these comparisons to uncover any relationship by: using a comprehensive interior structure model that thoroughly accounts for uncertainties in major element compositions (Mg, Si, Fe) and using a Bayesian statistical analysis which is more appropriate than other statistical methods used so far Hogg2010; Andreon2015 as well as properly accounting for the exoplanets at the rocky-to-volatile composition boundary. That is, we weigh each exoplanet proportional to the total probability of them being rocky given the uncertainties, thus including possibly volatile planets in a straightforward manner. This method is superior in comparisons to assuming negative CMF (non-physical) and more consistent than having assumptions about the total volatile content, which is a dimensionally degenerate problem.
@@
With the self-consistent data and our careful methodology we find that, although planets have a larger range in composition than their stars, there is an intriguing 1:1 relation. We investigate these discrepancies between our result and that of Adibekyan2021, Adibekyan2024 and Brinkman2024 (hereafter Adibekyan2021, Adibekyan2024 and Brinkman2024) by applying our methodology to their data and find that the main culprits are: the treatment of stellar data, choice of transit depth (more so than radial velocity) and the statistical modeling, such as consideration for the planets at the volatile rocky boundary. While a 1:1 relation is supportive of a primordial origin and thus compelling, the sample size and error bars are too large to make definite conclusions as a zero slope is also consistent with the data at the level. Therefore, we investigate how large of a sample of exoplanets is needed in order to firmly establish or rule out a direct connection between stars and their planets. It is important to note though, that although the data set we obtain is self consistent in terms of stellar data, it suffers from being observationally heterogeneous with transit depths and radial velocities being obtained using different instruments and retrieval codes.
2 Methods
2.1 Data selection
@@
In our previous study Plotnykov2020 we had chosen planets within the rocky M-R region (see Fig. 3) with good precision (). For this study we select exoplanets among that sample that orbit FGK stars with high-resolution optical spectra (R) and have high signal-to-noise ratios (S/N at Å) that enables host star chemical characterization with good precision. We also select for the most likely rocky exoplanets by checking that the equilibrium temperature is beyond the supercritical point of water () such that any volatile layer would be in gaseous form, vulnerable to atmospheric loss under high irradiation. We consider that their total mass and radius is consistent with a purely rocky composition. However, we acknowledge the possibility that some of these exoplanets may have a thick high molecular weight atmosphere, either CO/CO2 (’puffy Venuses’ applicable to K, Peng2024) or water (vapor planets applicable to K, Dorn2021) above a magma ocean so that their true rocky interior would be denser. Excluding these scenarios means our inferred Fe-MF is a minimum value. That is, planets can always have a larger proportion of iron to offset a volatile layer.
@@
Our sample consists of 21 exoplanets and 20 host stars, for which we derive homogeneous and physically self-consistent host star data of effective temperature, surface gravity, metallicity, masses and radii () (presented in paperI). Our methodology is described in Reggiani2022; Reggiani2024 and summarized here. We use the classical spectroscopy analysis to derive the initial set of photospheric stellar parameters (), which we than combine with GAIA astrometry and high-quality photometry data in our isochrone analysis (see paperI for more details). Both spectroscopic and photometric data is coupled together to derive the stellar parameters using a Markov Chain Monte Carlo (MCMC) framework. We iterate the fitting process multiple times to achieve self-consistency across all stellar parameters. This ensures self-consistency and robustness, reducing degeneracies and biases that could otherwise affect the accuracy of the abundances. Subsequently, we obtain individual elemental abundances ([Fe/H], [Mg/H], [Si/H], [Ca/H], [Al/H], [Ni/H]) directly from spectroscopy data using equivalent width-based calculation and manually verifying each absorption line feature, while excluding any fits that failed due to blending, line saturation or some other reason (see paperI).
@@
For comparison, Adibekyan2021/Adibekyan2024 used a similar method to derive their abundances, but lacked the isochrone analysis that updates and ensures consistency. On the other hand, Brinkman2024 uses a data-driven machine learning model (KeckSpec, a version of Cannon Ness2015) that uses Brewer2016; Brewer2018 data as the training set. Although convenient and fast, with this method it is difficult to assign individual uncertainties to predictions. We obtain similar results in terms of , but our physically self-consistent approach yields different elemental abundances with important consequences. We attribute this discrepancies to the different algorithms and selection processes employed. In addition, in contrast to all other studies, we apply non-local thermal equilibrium (non-LTE) corrections to all the elemental abundances, thereby including an important source of uncertainty. These corrections affect mostly the [Fe/H] and [Si/H] values, but not the [Mg/H].
@@
With self-consistent stellar data, we use radial velocity and transit measurements found in the NASA Exoplanet Archive (Akeson2013) to update planetary M-R of all planets. Of importance is that most of the exoplanets in our sample have been observed and analyzed by multiple groups and techniques (20 out of 21). We select those that have the largest number of measurements and highest precision, while making sure that transit observation comes from the same source instrument. Our choice for RV and transit data differs in some cases from those of Adibekyan2021, Adibekyan2024 and Brinkman2024 yielding another source for discrepancy in planetary M-R data (see Sec. 3.3 for more details).
2.2 Interior structure model
@@
To infer the planetary composition given planetary M-R, we use the interior structure model SuperEarth developed by Valencia2006; Valencia2007 and revamped by Plotnykov2020. SuperEarth code is a numerical interior integrator that solves for the mass, density, pressure, gravity and temperature as a function of radius, given a total mass, core-mass fraction (CMF) and chemical mineralogy (). That is, SuperEarth is a forward model that calculates the planet’s radius given its mass and interior structure: . In our analysis, we consider mineralogies based on the Earth and account for the following variations:
@@
To estimate interior composition from M-R data, we combine our forward interior model with an affine invariant Markov Chain Monte Carlo (MCMC) sampler EMCEE (Foreman2013), assuming that the M-R constraints are independent and follow a Gaussian distribution in our log-likelihood function. We check the effects of existing correlations between the stellar M-R posterior results (see paperI), but find that they do not translate to the planetary parameters, as the latter is dominated by uncertainties in the radial velocity/transit signal.
@@
For our MCMC setup, we follow the computational procedure outlined in Plotnykov2020 and assume uniform distributions within physical bounds for all the parameters (M, R, CMF and ), but for we specifically assume that by mol and by mol, while for the amount of wustite [(Mg,Fe)O], which affects the Mg/Si ratio, we adopt a fixed value (xWu). We have increased the xSi range compared to our previous study (Plotnykov2020) to encompass the wide range of proposed compositions consistent with Earth’s density profile (Fischer2015; Zhang2018; Umemoto2020; Hikosaka2022). Meanwhile, fixing xWu has minimal impact on the density profile of the planet and consequently on CMF ( (Plotnykov2024)).
@@
We had previously found that CMF values are highly dependent on prior assumptions (e.g. xSi, xFe)222The change in CMF can be up to ., while to the iron-mass fraction (Fe-MF) of planets remains highly invariant to most mineralogy priors and where variations do occur, they tend to be minor Plotnykov2024. Thus, for this study we focus on Fe-MF values and obtain the full range by considering, a-posteriori, a mineralogical model consistent with the density profile that fits the planetary M-R. For this, we consider minor minerals such as corundum () and calcium perovskite () in the mantle because these are major carriers of Al and Ca (xAl, xCa) as well as we vary nickle in the core for the same reason (xNi). In contrast to paperI, we also vary the Mg/Si ratio controlled by the proportions of mantle minerals (e.g. bridgmanite vs wustite). Note that in the case when the mantle Mg/Si molar ratio is less than 1, in the absence of experiments, we assume that there is no wustite in the mantle (xWu=0) and the remaining silica is incorporated into quartz (, xQtz). However, experiments should be carried out to better understand the mineralogy and equation of state (EOS) at these low Mg/Si values, as there is a considerable number of stars that display them ( of APOGEE stars have Mg/Si by weight APOGEE).
@@
Thus, owing to the lack of geophysical and theoretical constraints in these relative abundances, we adopt host star elemental ratios by weight for Mg/Si, Ca/Mg, Al/Mg and Ni/Fe as priors to constrain the added mineralogy. Almost all of these ratios are consistent among stars with different metallicities. The exception is the Mg/Si ratio, which has the most variation among the different stars when compared to GALAH (GALAH), APOGEE (APOGEE), HYPATIA (HYPATIA) surveys as well as the variation within the solar system (Anders1964).
@@
Adding these minor minerals a-posteriori has a negligible effect in the density and CMF, given that their EOS’s are fairly similar to the major chemical phases our interior structure model considers333The difference between using the minor mineralogies for calculating the structure versus after is a difference in radius of km for a 1 earth-mass planet.. However, they are needed to obtain the true Fe-MF range allowed for a planet consistent with its M-R. For example, depending on the amount of these minor elements, the Fe-MF can change by (Plotnykov2024).
@@
2.3 Stellar equivalent chemical model
@@
2.4 Statistical method
Owing to the small sample size (N=21), we choose to investigate possible correlations between stellar and planetary parameters with a simple linear function () as to avoid over-fitting the data. To properly evaluate the best-fit line, we use STAN codebase (stan) and set-up a Hamiltonian Monte Carlo (HMC) with a No-U-Turn sampler (NUTS) to find the posterior for (gradient) and (intercept) values. This method is superior in comparison to ordinary least-square (OLS) or orthogonal linear regression (ODR), due to a multitude of factors Hogg2010; Andreon2015. Specifically, in the context of our sample, a Bayesian method mitigates the effect of outliers and allows for fitted parameters to have non-Gaussian posterior distributions ( and ), both of which are extremely important factors owing to the small sample size (see Sec. 3.3 for comparison). More importantly, however, is that a Bayesian analysis has greater flexibility on the assumed probability density of the input data, in contrast to both OLS and ODR treatments that assume symmetrical errors (Gaussian distribution). We find that most stellar and planetary posterior parameters (e.g. Fe-MFst, Fe-MFpl) have skewness in their distributions, thereby we assume each parameter follows a skewed-Gaussian () distribution in our likelihood () analysis.

### v2

@@ I Introduction
Starting with planetesimals, that are presumed to have homogenous composition (similar to the refractory ratios of the host star Larimer 1967; Grossman 1972) under efficient radial mixing in the protoplanetary disk, the planet grows through a series of processes that may or may not alter their primordial composition (Asphaug et al., 2006; Carter et al., 2015; Scora et al., 2020; Clement et al., 2021). In the Solar System, there is a moderate variety in major element composition as seen by the iron enrichment of Mercury and stony-iron meteorites as well as the iron-depletion of the Moon and aubrite meteorites, with values in between (Keil, 1989; Morgan and Anders, 1980; Lodders, 2003; Wieczorek, 2006; Margot et al., 2018). Proposed mechanisms that result in major chemical sorting so far include those that are based on differing Fe, Si and Mg properties early in the evolution, such as, fractionation driven by the differences in the condensation temperatures of Fe vs Mg-Si (Lewis, 1972; Grossman, 1972; Aguichine et al., 2020); photophoresis, which causes Fe vs Mg-Si dust grains to drift differently due to differences in conductivity and density when heated by stellar radiation (Krauss and Wurm, 2005; Wurm et al., 2013; Cuello et al., 2016); increased nucleation of Fe vs Mg-Si pebbles due to differences in grain size and surface tension that affect growth under the streaming instability (Johansen and Dorn, 2022); or those that are based on losing mantle material once the planet has differentiated into a silicate mantle over an iron core, such as giant impact collisions (Benz et al., 1988, 2007; Asphaug and Reufer, 2014; Clement et al., 2021; Scora et al., 2024; Ferich et al., 2025) and for small bodies (i.e. Mercury-sized), photoevaporation of the mantle (Ito and Ikoma, 2021). Thus, comparing the relative abundances of planetary bodies to that of their star, offers the possibility to understand which mechanisms form planets.
@@
Chemical comparisons had been limited only to the Solar System or a few planets (Santos et al., 2015) until recently, when enough data of rocky exoplanets with good precision had grown to several tens of planets. Plotnykov and Valencia (2020) compared a population of allegedly rocky planets to that of the population of planet-hosting stars and found that planets exhibit a greater diversity in core-mass fraction (CMF), iron to magnesium (Fe/Mg) and iron to silicate (Fe/Si) ratios relative to the stellar population. Subsequent studies made one-to-one comparisons in an effort to uncover any possible correlations (Adibekyan et al., 2021; Schulze et al., 2021; Wang et al., 2022; Liu and Ni, 2023; Unterborn et al., 2023; Brinkman et al., 2024; Adibekyan et al., 2024). Notably, Adibekyan et al. (2021) suggested that super-Mercuries form a distinct population around stars that are enriched in core to mantle material (Fe/Mg+Si) and proposed a strong linear relationship with a slope steeper than a 1:1 line between iron-to-silicate mass fraction of planets and their host stars (slope of ). This result has been revised in Brinkman et al. (2024) and Adibekyan et al. (2024) works such that a super-Mercury subpopulation is no longer present, but the steep compositional slope remains.
@@
However, most of these different studies have a shortcoming that stems from considering separately the elemental abundances of stars versus their stellar mass and radius. Given that the latter affect inferences on the planetary mass-radius (M-R) and consequently the planetary interior composition, ensuring consistency between stellar composition, mass and radii is expected to update the planetary characteristics. To address this issue and ensure consistency, two groups (Brinkman et al., 2024; Ross et al., 2025) have reanalyzed the stellar data and updated planetary M-R accordingly. In our companion paper, hereafter paper I (Ross et al., 2025), we explain in detail the procedure to obtain a consistent stellar data set and perform simple planet-stars comparisons.
@@
In this paper, we delve deeper into these comparisons to uncover any relationship by: using a comprehensive interior structure model that thoroughly accounts for uncertainties in major element compositions (Mg, Si, Fe) and using a Bayesian statistical analysis which is more appropriate than other statistical methods used so far Hogg et al. (2010); Andreon and Weaver (2015) as well as properly accounting for the exoplanets at the rocky-to-volatile composition boundary. That is, we weigh each exoplanet proportional to the total probability of them being rocky given the uncertainties, thus including possibly volatile planets in a straightforward manner. This method is superior in comparisons to assuming negative CMF (non-physical) and more consistent than having assumptions about the total volatile content, which is a dimensionally degenerate problem.
@@
With the self-consistent data and our careful methodology we find that, although planets have a larger range in composition than their stars, there is an intriguing 1:1 relation. We investigate these discrepancies between our result and that of Adibekyan et al. (2021), Adibekyan et al. (2024) and Brinkman et al. (2024) (hereafter A21, A24 and B24) by applying our methodology to their data and find that the main culprits are: the treatment of stellar data, choice of transit depth (more so than radial velocity) and the statistical modeling, such as consideration for the planets at the volatile rocky boundary. While a 1:1 relation is supportive of a primordial origin and thus compelling, the sample size and error bars are too large to make definite conclusions as a zero slope is also consistent with the data at the level. Therefore, we investigate how large of a sample of exoplanets is needed in order to firmly establish or rule out a direct connection between stars and their planets. It is important to note though, that although the data set we obtain is self consistent in terms of stellar data, it suffers from being observationally heterogeneous with transit depths and radial velocities being obtained using different instruments and retrieval codes.
II Methods
II.1 Data selection
@@
In our previous study Plotnykov and Valencia (2020) we had chosen planets within the rocky M-R region (see Fig. 3) with good precision (). For this study we select exoplanets among that sample that orbit FGK stars with high-resolution optical spectra (R) and have high signal-to-noise ratios (S/N at Å) that enables host star chemical characterization with good precision. We also select for the most likely rocky exoplanets by checking that the equilibrium temperature is beyond the supercritical point of water () such that any volatile layer would be in gaseous form, vulnerable to atmospheric loss under high irradiation. We consider that their total mass and radius is consistent with a purely rocky composition. However, we acknowledge the possibility that some of these exoplanets may have a thick high molecular weight atmosphere, either CO/CO2 (’puffy Venuses’ applicable to K, Peng and Valencia (2024)) or water (vapor planets applicable to K, Dorn and Lichtenberg (2021)) above a magma ocean so that their true rocky interior would be denser. Excluding these scenarios means our inferred Fe-MF is a minimum value. That is, planets can always have a larger proportion of iron to offset a volatile layer.
@@
Our sample consists of 21 exoplanets and 20 host stars, for which we derive homogeneous and physically self-consistent host star data of effective temperature, surface gravity, metallicity, masses and radii () (presented in paper I). Our methodology is described in Reggiani et al. (2022, 2024) and summarized here. We use the classical spectroscopy analysis to derive the initial set of photospheric stellar parameters (), which we than combine with GAIA astrometry and high-quality photometry data in our isochrone analysis (see paper I for more details). Both spectroscopic and photometric data is coupled together to derive the stellar parameters using a Markov Chain Monte Carlo (MCMC) framework. We iterate the fitting process multiple times to achieve self-consistency across all stellar parameters. This ensures self-consistency and robustness, reducing degeneracies and biases that could otherwise affect the accuracy of the abundances. Subsequently, we obtain individual elemental abundances ([Fe/H], [Mg/H], [Si/H], [Ca/H], [Al/H], [Ni/H]) directly from spectroscopy data using equivalent width-based calculation and manually verifying each absorption line feature, while excluding any fits that failed due to blending, line saturation or some other reason (see paper I).
@@
For comparison, A21/A24 used a similar method to derive their abundances, but lacked the isochrone analysis that updates and ensures consistency. On the other hand, B24 uses a data-driven machine learning model (KeckSpec, a version of Cannon Ness et al. (2015)) that uses Brewer et al. (2016); Brewer and Fischer (2018) data as the training set. Although convenient and fast, with this method it is difficult to assign individual uncertainties to predictions. We obtain similar results in terms of , but our physically self-consistent approach yields different elemental abundances with important consequences. We attribute this discrepancies to the different algorithms and selection processes employed. In addition, in contrast to all other studies, we apply non-local thermal equilibrium (non-LTE) corrections to all the elemental abundances, thereby including an important source of uncertainty. These corrections affect mostly the [Fe/H] and [Si/H] values, but not the [Mg/H].
@@
With self-consistent stellar data, we use radial velocity and transit measurements found in the NASA Exoplanet Archive (Akeson et al., 2013) to update planetary M-R of all planets. Of importance is that most of the exoplanets in our sample have been observed and analyzed by multiple groups and techniques (20 out of 21). We select those that have the largest number of measurements and highest precision, while making sure that transit observation comes from the same source instrument. Our choice for RV and transit data differs in some cases from those of A21, A24 and B24 yielding another source for discrepancy in planetary M-R data (see Sec. III.3 for more details).
II.2 Interior structure model
@@
To infer the planetary composition given planetary M-R, we use the interior structure model SuperEarth developed by Valencia et al. 2006, 2007 and revamped by Plotnykov and Valencia (2020). SuperEarth code is a numerical interior integrator that solves for the mass, density, pressure, gravity and temperature as a function of radius, given a total mass, core-mass fraction (CMF) and chemical mineralogy (). That is, SuperEarth is a forward model that calculates the planet’s radius given its mass and interior structure: . In our analysis, we consider mineralogies based on the Earth and account for the following variations:
@@
To estimate interior composition from M-R data, we combine our forward interior model with an affine invariant Markov Chain Monte Carlo (MCMC) sampler EMCEE (Foreman-Mackey et al., 2013), assuming that the M-R constraints are independent and follow a Gaussian distribution in our log-likelihood function. We check the effects of existing correlations between the stellar M-R posterior results (see paper I), but find that they do not translate to the planetary parameters, as the latter is dominated by uncertainties in the radial velocity/transit signal.
@@
For our MCMC setup, we follow the computational procedure outlined in Plotnykov and Valencia (2020) and assume uniform distributions within physical bounds for all the parameters (M, R, CMF and ), but for we specifically assume that by mol and by mol, while for the amount of wustite [(Mg,Fe)O], which affects the Mg/Si ratio, we adopt a fixed value (xWu). We have increased the xSi range compared to our previous study (Plotnykov and Valencia, 2020) to encompass the wide range of proposed compositions consistent with Earth’s density profile (Fischer et al., 2015; Zhang et al., 2018; Umemoto and Hirose, 2020; Hikosaka et al., 2022). Meanwhile, fixing xWu has minimal impact on the density profile of the planet and consequently on CMF ( (Plotnykov and Valencia, 2024)).
@@
We had previously found that CMF values are highly dependent on prior assumptions (e.g. xSi, xFe)222The change in CMF can be up to ., while to the iron-mass fraction (Fe-MF) of planets remains highly invariant to most mineralogy priors and where variations do occur, they tend to be minor Plotnykov and Valencia (2024). Thus, for this study we focus on Fe-MF values and obtain the full range by considering, a-posteriori, a mineralogical model consistent with the density profile that fits the planetary M-R. For this, we consider minor minerals such as corundum () and calcium perovskite () in the mantle because these are major carriers of Al and Ca (xAl, xCa) as well as we vary nickle in the core for the same reason (xNi). In contrast to paper I, we also vary the Mg/Si ratio controlled by the proportions of mantle minerals (e.g. bridgmanite vs wustite). Note that in the case when the mantle Mg/Si molar ratio is less than 1, in the absence of experiments, we assume that there is no wustite in the mantle (xWu=0) and the remaining silica is incorporated into quartz (, xQtz). However, experiments should be carried out to better understand the mineralogy and equation of state (EOS) at these low Mg/Si values, as there is a considerable number of stars that display them ( of APOGEE stars have Mg/Si by weight Majewski et al. (2017)).
@@
Thus, owing to the lack of geophysical and theoretical constraints in these relative abundances, we adopt host star elemental ratios by weight for Mg/Si, Ca/Mg, Al/Mg and Ni/Fe as priors to constrain the added mineralogy. Almost all of these ratios are consistent among stars with different metallicities. The exception is the Mg/Si ratio, which has the most variation among the different stars when compared to GALAH (Buder et al., 2024), APOGEE (Majewski et al., 2017), HYPATIA (Hinkel et al., 2014) surveys as well as the variation within the solar system (Anders, 1964).
@@
Adding these minor minerals a-posteriori has a negligible effect in the density and CMF, given that their EOS’s are fairly similar to the major chemical phases our interior structure model considers333The difference between using the minor mineralogies for calculating the structure versus after is a difference in radius of km for a 1 earth-mass planet.. However, they are needed to obtain the true Fe-MF range allowed for a planet consistent with its M-R. For example, depending on the amount of these minor elements, the Fe-MF can change by (Plotnykov and Valencia, 2024).
@@
II.3 Stellar equivalent chemical model
@@
II.4 Statistical method
Owing to the small sample size (N=21), we choose to investigate possible correlations between stellar and planetary parameters with a simple linear function () as to avoid over-fitting the data. To properly evaluate the best-fit line, we use STAN codebase (Stan Development Team, 2024) and set-up a Hamiltonian Monte Carlo (HMC) with a No-U-Turn sampler (NUTS) to find the posterior for (gradient) and (intercept) values. This method is superior in comparison to ordinary least-square (OLS) or orthogonal linear regression (ODR), due to a multitude of factors Hogg et al. (2010); Andreon and Weaver (2015). Specifically, in the context of our sample, a Bayesian method mitigates the effect of outliers and allows for fitted parameters to have non-Gaussian posterior distributions ( and ), both of which are extremely important factors owing to the small sample size (see Sec. III.3 for comparison). More importantly, however, is that a Bayesian analysis has greater flexibility on the assumed probability density of the input data, in contrast to both OLS and ODR treatments that assume symmetrical errors (Gaussian distribution). We find that most stellar and planetary posterior parameters (e.g. Fe-MFst, Fe-MFpl) have skewness in their distributions, thereby we assume each parameter follows a skewed-Gaussian () distribution in our likelihood () analysis.